OpenStudy (anonymous):

I REALLY REALLY NEED HELP! Will give medal for right answer! Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4^x and y = 2^x + 2 intersect are the solutions of the equation 4^x = 2^x + 2. Part B: Make tables to find the solution to 4^x = 2^x + 2. Take the integer values of x between -3 and 3. Part C: How can you solve the equation 4^x = 2^(x + 2) graphically? (Note: when it says 2^x + 2 it means 2 raised to the power of x+2. x+2 is the exponent)

OpenStudy (amistre64):

when the graphs of 2 functions meet, what does that tell us about the values of x,y ?

OpenStudy (anonymous):

It means that (x, y) is the solution?

OpenStudy (amistre64):

yes, but why would only 1 point be a solution? what property do the graphs share when they are at the same point?

OpenStudy (amistre64):

y = 4^x and y = 2^x + 2 intersect at a common point when they are equal to each other .....

OpenStudy (amistre64):

if they do not have a point in common, they are NOT equal to each other .....

OpenStudy (amistre64):